OpenStudy (anonymous):

if a triangle has an area of 27cm and a base length of 9, calculate the height of the triangle to that base.

OpenStudy (ghazi):

area of triangle = (1/2)*base * height 27=(1/2)*9*base

OpenStudy (anonymous):

ok

OpenStudy (ghazi):

@oholmes can you do this?

OpenStudy (anonymous):

6?

OpenStudy (ghazi):

yes